• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 00:06:26 on October15,2009

    Entry number 294968

    keyword=Compton run summary

    Run 20399: electron only run at 7.0mm, 2 plane coinc, FADC trig on edet, 
    electron prescale 100%, to check background.  Rates very very low (8Hz/uA laser 
    Run 20400: electron only run moving around the electron detector.  Everything 
    else the same.
    Run 20401: electron only run moving around the electron detector.  Rates are no 
    good at -3.0mm.  -2.0mm rates still not okay (no difference between laser 
    locked and unlocked).  Okay at -1.0.
    Run 20402: edet at -1.0mm.  Width strips 2.  Otherwise the same.
    Run 20403: e-g coincidence run still at -1.0mm, standard FADC trigger, 2 plane 
    coincidence, prescales 4%, 100%, 100%, width strips still at 2 (95uA).
    Run 20404: same as last run, width strips 5.  
    Run 20406: photon rate check (coincidence run), otherwise same as last run. 
    (prescales 4%, 100%, 100%, 2 planes, width strips 5). 15uA (with closing slit). 
     ~10 minutes.
    Run 20407: 15uA coincidence run, prescales 4%, 0%, 100%.